OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2-2x= 99

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

You still there for me to answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok. So you will get x^2 - 2x - 99 = 0

OpenStudy (anonymous):

\[x^2-2x= 99 \\ (x-1)^2=99+1=100\\ x-1=10,x-1=-10\\ x=11,x=-9\]

OpenStudy (anonymous):

@satellite73 That is not solving using the quadratic formula.

OpenStudy (anonymous):

It has to be quadratic formula.

OpenStudy (anonymous):

You use the coefficient of X^2 - 2x - 99 = 0 in order to find the coefficients and thus determine a, b and c.

OpenStudy (anonymous):

So, a = 1, b = -2 and c = -99

OpenStudy (anonymous):

A= 1 B= -2 C= -99

OpenStudy (anonymous):

I get 20 as the discriminant

OpenStudy (anonymous):

\[(-b \pm \sqrt{b ^{2}-4ac})\div2a\]

OpenStudy (anonymous):

So you plug the numbers given above into the quadratic formula to come up with 2 solutions.

OpenStudy (anonymous):

I don't get 11 and -9

OpenStudy (anonymous):

Don't think that is what you're meant to get. I'll try it.

OpenStudy (anonymous):

The answer is -9 and 11. I just don't know how to get it.

OpenStudy (anonymous):

Is it possible that this includes an imaginary number?

OpenStudy (anonymous):

Oh no never mind.

OpenStudy (anonymous):

Yes I got the correct result using the quadratic formula.

OpenStudy (anonymous):

Don't forget that subtracting a negative makes it a positive.

OpenStudy (anonymous):

Can you show me :x

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{4-(4-99)} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{400} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 2\pm20 }{ 2 }\]

OpenStudy (anonymous):

\[1\pm10\]

OpenStudy (anonymous):

So 1 + 10 = 11 and 1 - 10 = -9.

OpenStudy (anonymous):

I'm confused on how you got 4-(4-99)

OpenStudy (anonymous):

Sorry. That was meant to be 4*-99. And the 4 you get by doing b^2, so -2^2.